### 13 Apr 2026 v2

# Modular memristor model with synaptic-like plasticity and volatile memory

Abstract
Compact models of memristors are essential for simulating large-scale neuromorphic systems, yet they often do not include description of complex dynamics like volatile relaxation and synaptic plasticity. We introduce a modular, computationally efficient memristor model that bridges this gap by integrating principles from physics and computational neuroscience. Starting from standard memristive system dynamics, the framework incorporates synaptic-like plasticity dynamics, a mapping from state variables to cumulative conductance, a volatility module, and a saturation module. The plasticity component is inspired by a biological rule for spike-timing-dependent plasticity (STDP) and is compatible with the general memristive systems formalism. Finally, we propose a Laplace transform-based technique to derive the precise form of the mapping from state variables to cumulative conductance, replacing ad hoc voltage-current relationships with principled construction.
We quantitatively validate the complete model against a comprehensive set of experimental data from polymeric memristors exhibiting potentiation, synaptic-like plasticity and volatile decay. Our work presents a new paradigm for memristor modeling that is both practical for large-scale simulation and rich in explanatory power, providing a principled tool for the design of next-generation neuromorphic hardware.
Teaser: Kernel-based volatility and eligibility-trace plasticity unify short-term memory and learning in polymer memristor compact models.
Introduction
Neuromorphic computing allows physical devices to bring memory and computation closer together, targeting improvements in energy efficiency compared to traditional von Neumann architectures [28, 9]. Memristors are one-port devices whose conductance depends on the history of electrical stimulation [2, 1] and are particularly well suited to serve as artificial synapses [10]. Their inherent memory, analog tunability, and scalability make them attractive for both fundamental studies and applications in neuromorphic hardware [28].
Compact models such as TEAM and VTEAM [13, 14] provide versatile descriptions of memristive switching [27], widely used in circuit-level simulations. However additional properties must be accounted for in a quantitative and flexible manner for applications in neural networks, such as conductance volatility supporting short-term memory and fading dynamics [26, 11, 25] and synaptic-like plasticity enabling learning and long-term adaptation [10, 29, 20]. The recently proposed V-VTEAM model [18] addresses volatility by modifying the dynamics of the memristive system, but a unified framework that combines synaptic-like plasticity, volatility and saturation while remaining modular and computationally efficient is still lacking.
We propose a modular memristor model that augments the standard memristive system description with modules for plasticity, volatility and saturation. This framework unifies:
  1. 1.
    a standard memristive systems component utilizing established voltage-driven models [23, 14];
  2. 2.
    a novel synaptic-like plasticity component based on local-variable eligibility traces [6], yielding biologically inspired causal and anti-causal weight updates;
  3. 3.
    a cumulative conductance function that maps state variables to conductance;
  4. 4.
    a volatility module inspired by linear viscoelasticity [3, 15], where material response depends on a hereditary convolution kernel; and
  5. 5.
    a saturation module implemented as a linear-nonlinear scheme in conjunction with volatility, to produce bounded device conductance.
One can select the memristive core, cumulative conductance mapping, decay kernel and saturation function to match a given device and enable or disable volatility and/or STDP without changing the underlying modeling framework. Although we present the framework in voltage-driven form consistent with common experimental protocols, its modular structure also permits straightforward translation to current-driven variants.
We validate the approach using experimental data from polymeric memristors developed by Pfleger, Foulger, and co-workers [17, 16, 8, 5]. These memristors consist of a thin film of poly(N-(3-(9H-carbazol-9-yl)propyl)methacrylamide) sandwiched between ITO and top metallic electrodes. The devices studied exhibited pronounced hysteresis in their I-V characteristics, bistable resistive switching effects at higher voltages of +/-5 V (with a maximum ON/OFF ratio of around 200 and retention time exceeding 2 hours)[17], and analog, pulse-driven conductance modulation at lower voltages that resembles potentiation and depression of neuronal synapses [16]. For this class of devices, the internal conductance state depends on the history of applied voltage via a convolution with a kernel that decays approximately as $1/t$. This form reflects a broad distribution of relaxation times, consistent with disordered systems and percolation phenomena in which conduction pathways evolve stochastically over time [22, 21].
Using an appropriate measurement protocol, the distinct effects of the individual components may be effectively separated. To this end, we developed a measurement procedure that exploits the modular structure of the framework to efficiently estimate the model parameters, enabling near-independent optimization of each functional component.
Finally, we introduce a technique based on the Laplace transform to determine the functional form of the cumulative conductance mapping for a given device. Owing to the simplicity of the model framework, this form may provide useful insight to the characteristic macroscopic behavior of the device.
The model reproduces both volatile memory and synaptic plasticity observed in polymeric memristors, while maintaining computational efficiency for large-scale neuromorphic simulations. We believe this approach will help bridge the gap between device-level physics and system-level neural computation, with particular relevance for sustainable, polymer-based memristive hardware.
Results
We present the theoretical model and its application to hardware for parameter fitting.
Model
For reference, we recall the general framework for voltage-controlled, time-invariant one-port memristive systems introduced by Chua and Kang [1]. Such a device is described by a differential equation governing the evolution of a vector of internal state variables $\bm{x}$, together with an algebraic equation relating the current and voltage through a conductance function $G(\bm{x})$:
$i$ $=G(\bm{x})\,v$ (1)
$\dot{\bm{x}}$ $=f(\bm{x},v)\,.$
[…]
Our approach is inspired by the memristive system framework, but differs from it in several important aspects. Following Eq.˜1, we introduce a vector of four internal state variables: s,w,x,ys,\,w,\,x,\,y, along with a description of their respective dynamics. To incorporate STDP into our memristor model, we introduce a formulation of spike timing-dependent plasticity dynamics (variables w,xw,\,x and $y$) for one-port electronic devices, compatible with the standard formulation of memristive systems.
However, unlike in the formulation above, our device conductance function $G$ is not directly a function of the state parameters. Instead, we introduce a kernel function $\ker(\,)$ to capture the temporal decay of conductance, a "cumulative conductance" mapping $H(s,w)$ from state variables to non-volatile device conductance and a saturation function $\operatorname{sat}(\,)$ to model device conductance saturation. A schematic of the modules’ relationships is provided in Figure˜1.
Figure 1: Schematic representation of the model. The model combines two independent components into the cumulative conductance $H(s,w)$: voltage-controlled switching dynamics providing the memristive core (state variable $s$) , and STDP-like synaptic plasticity implementing learning rules via eligibility traces (state variable $w$). The cumulative conductance then undergoes volatile decay through viscoelastic-inspired convolution with a decay kernel $\ker()$ and the resulting device conductance $G$ emerges after a saturation nonlinearity.
Memristive system dynamics
This component builds upon established voltage-driven memristive models, suitable as the basis for non-volatile behavior. A simple and common possibility for the memristive model core is the linear voltage controlled memristor model [23]:
$\dot{s}=\mu\,v\,.$ (2)
A more detailed approach due to [14] is also a popular choice.
Notably, we have omitted the windowing functions in the above equation, because in our approach the state parameters are unbounded at this stage of the modeling process. Hence the so-called linear memristor model of Eq.˜2 is truly a linear differential equation at this stage. We resolve non-physical behavior by the saturation component of the model.
It should be emphasized that any kind of voltage-controlled memristor state-variable differential equation in the form specified in Eq.˜1 is compatible with our framework.
Adaptation of STDP to a memristive device
In this section, we adapt the local variables synaptic STDP model, as presented in [6], to a one-port physical electronic device, making it compatible with the standard memristive systems framework. The model provides a phenomenological description of synaptic weight parameter changes based on relative timing of two classes of spiking inputs: presynaptic and postsynaptic. Notably, due to its abstract formulation, it can be generalized beyond neural systems, provided the two input classes are reinterpreted in an appropriate manner.
To begin, recall the synaptic model’s original formulation and interpretation. We consider the synaptic weight variable $w$, whose dynamics are driven by presynaptic and postsynaptic spike trains $S_{\mathrm{pre}}$ and $S_{\mathrm{post}}$, respectively. Denote by $(t_{f})$ and $(t^{n})$ the presynaptic and postsynaptic spike time sequences. The spike trains are then modeled as a series of Dirac impulses at the corresponding times:
$S_{\mathrm{pre}}(t)=\sum_{f}\delta(t-t_{f}),$ $\text{and}\quad S_{\mathrm{post}}(t)=\sum_{n}\delta(t-t^{n}).$
Define two auxiliary variables $x$ and $y$, representing the presynaptic and postsynaptic traces, which model the capacity for change in the weight variable $w$. They are incremented by their respective spike trains and decay exponentially between spikes:
$\dot{x}(t)$ $=-\frac{x(t)}{\tau_{\mathrm{+}}}+S_{\mathrm{pre}}(t)$
$\dot{y}(t)$ $=-\frac{y(t)}{\tau_{\mathrm{-}}}+S_{\mathrm{post}}(t),$
where $\tau_{\mathrm{+}},\,\tau_{\mathrm{-}}$ are time constants.
Finally, the weight variable $w$ is updated at postsynaptic spike times proportionally to the presynaptic trace $x$, and negatively at postsynaptic spike times in proportion to the postsynaptic trace $y$:
$\dot{w}(t)$ $=x(t)\,A_{\mathrm{+}}(w)\,S_{\mathrm{post}}(t)-y(t)\,A_{\mathrm{-}}(w)\,S_{\mathrm{pre}}(t)\,.$
The functions $A_{\mathrm{+}},\,A_{\mathrm{-}}$ are weight dependent scaling factors.
Our aim is to reformulate this model for a one-port device, which differs from biological synapses in several important respects. The key idea is to replace discrete spike trains with continuous voltage signals, while interpreting one terminal as the ’presynaptic’ terminal and the other as the ’postsynaptic’ terminal. With this identification, the respective spike trains correspond to positive and negative voltages across the device. Accordingly, we substitute the spike trains at each synaptic terminal with the appropriate signed voltage function, as illustrated below:
$S_{\mathrm{pre}}(t)$ $\implies v^{+}(t)$
$S_{\mathrm{post}}(t)$ $\implies v^{-}(t)\,$,
where $v^{+},\,v^{-}$ are the positive and negative parts of $v$, respectively, with the convention $v=v^{+}-v^{-}$.
Adding variables w,x,yw,\,x,\ y as additional state parameters of the memristive device model, we can formulate the synaptic-like plasticity component for the one-port device as follows:
$\dot{x}$ $=-\frac{x}{\tau_{\mathrm{+}}}+v^{+}$ (3)
$\dot{y}$ $=-\frac{y}{\tau_{\mathrm{-}}}+v^{-}$
$\dot{w}$ $=x\,A_{\mathrm{+}}(w)\,v^{-}-y\,A_{\mathrm{-}}(w)\,v^{+}\,$,
where $\tau_{\mathrm{+}},\,\tau_{\mathrm{-}}$ are time constants and the functions $A_{\mathrm{+}},\,A_{\mathrm{-}}$ are weight dependent scaling factors. We denote by $v^{+},\,v^{-}$ the positive and negative parts of $v$, respectively, with the convention $v=v^{+}-v^{-}$.
We differ from [6] in our approach by considering unbounded $w$ at this stage. As a result the functions $A_{+},\,A_{-}$ do not need to contain a windowing function to force bounds onto the weight parameter. We note that in the model, the memristive system components may incorporate coupling by allowing $A_{\mathrm{+}}$ and $A_{\mathrm{-}}$ to depend on $s$.
Volatility kernel and cumulative conductance
In order to model memristor volatility characteristics, we implement a standard convolution-based approach, inspired by linear viscoelasticity theory [3]. A key advantage is that this approach enables the use of $1/t^{\alpha}$ volatility kernels that are otherwise computationally unavailable. We select a suitable kernel function $\ker()$, which characterizes the device’s conductance volatility, and the cumulative conductance function $H(s,w)$, which describes nonvolatile behavior of conductance. The resulting volatile cumulative conductance $H_{\mathrm{vol}}$ is calculated as the convolution
$H_{\mathrm{vol}}=\ker\ast\,dH(s,w).$ (4)
Note that the convolution is calculated with respect to time, since the model state variables $s$ and $w$ are time-dependent. Similarly, the differential $dH$ must also be taken with respect to time, written explicitly:
$dH=\partial_{s}H(s,w)\,\dot{s}+\partial_{w}H(s,w)\,\dot{w}\,.$ (5)
Regarding the kernel function, we may distinguish two cases, based on its limiting behavior. For example, if the kernel is the Heaviside function, no volatile behavior is present (pure “elastic” regime), and Eq.˜4 simplifies to
$H_{\mathrm{vol}}=H(s,w).$
This relation motivates the label "cumulative conductance" for the function $H$; it captures the functional relationship of state variables and device conductance in the non-volatile case. Conversely, if $\lim_{t\rightarrow\infty}\ker(t)=0$, we are in the viscoelastic regime, where the system decays to its equilibrium over time. Our model is therefore sufficiently general to describe devices exhibiting a combination of volatile and non-volatile characteristics.
The form of the kernel and cumulative conductance functions is device-specific and also depends on the chosen voltage-controlled model. Inspired by the presentation in Kvatinsky et al. [14], we provide ad-hoc variants for the kernel function and for the cumulative conductance function.
First, a linear and exponential relationship of the cumulative conductance on the state variables are presented:
$dH(s,w)$ $=h_{1}\,\dot{s}+h_{0}\,\dot{w}$
$dH(s,w)$ $=h_{1}\,e^{a\,s}\,\dot{s}+h_{0}\,e^{b\,w}\,\dot{w}\,.$
The terminology corresponds to the usual relationships in the case of the purely elastic regime ($\ker()$ is the Heaviside function), as discussed above. Thus both versions generalize the current-voltage relationships presented in Kvatinsky et al. [14].
Similarly, one can consider variants for the kernel:
$\ker(t)$ $=e^{-at}$
$\ker(t)$ $=\frac{1}{(t+\varepsilon)^{a}}\,$,
for $t>0$ and identically zero for negative times. The power-law kernel may be offset by a small positive $\varepsilon$ to avoid the singularity at zero.
In addition, we offer a rigorous but data-specific approach for deriving the functional forms of the volatility characteristics $\ker()$ and $H$ in Eq.˜9 and Eq.˜11, respectively. We do so using data from pulse stimulation of the device, idealized by a single box function. Consequently, in the fitting section below we consider a conductance function of the form
$dH(s,w)=(h_{11}\,|s|^{a}+h_{10})\dot{s}+(h_{01}|w|^{a}+h_{00})\dot{w}\,.$ (6)
As a final remark, we note that convolving the kernel with the derivative $dH$ is useful because it allows the convolution term to be effectively ’reset’ by the state parameter derivatives $\dot{s},\,\dot{w}$, which is necessary for the kernel-based approach to work correctly.
Saturation and resulting conductance
Since the approach described so far allows memristor conductance values that are unbounded and could even become negative, it is necessary to limit the conductance to realistic bounds. To obtain the resulting conductance from the volatile, unbounded version Eq.˜4, we employ the linear-nonlinear scheme [6]; a standard method to introduce nonlinearities, such as saturation, into a linear system.
Hence, we calculate the resulting conductance $G$ as:
$G=\operatorname{sat}(H_{\mathrm{vol}})\,$, (7)
where $\operatorname{sat}(\,)$ is a chosen nonlinear saturation function.
Sigmoid functions are a common choice for saturation functions given their flexibility, hence here we adopt the logistic function:
$\operatorname{sat}(h)=(g_{\mathrm{max}}-g_{\mathrm{min}})\frac{1}{1+e^{-A\,h}}+g_{\mathrm{min}}\,$, (8)
where
$A=\frac{4}{h_{\mathrm{max}}-h_{\mathrm{min}}}$
to best align the linear portion of the logistic function with the range of conductance.
Fitting to experimental data
Application of our model to experiments relies on comprehensive electrical characterization data, using protocols designed to isolate different aspects of memristive behavior. Measurements were done using trains of stimulation pulses (±500 mV, 20 ms duration) with interleaved read pulses (-50 mV, 50 ms duration) to monitor conductance evolution. Owing to observed voltage-threshold behavior, the contribution of the read pulses was considered negligible and therefore omitted from the model simulations.
Two measurement protocols formed the empirical basis for model development. First, relaxation measurements following stimulation at different intensities revealed volatile memory dynamics, with conductance monitored over timescales from tens of milliseconds to hundreds of seconds. Second, specialized timing protocols captured STDP characteristics by varying the delay between paired pulses, mapping the full spectrum of causal and anti-causal plasticity.
To derive the model for a given device, we first extract the characteristic volatility kernel from the relaxation measurements using an approximation technique. We then apply the Laplace method described below to determine the form of the cumulative function and estimate the corresponding parameters. Next, we estimate the parameters of the STDP module from measurements and finally set the saturation module coefficients either to minimally affect the model, or by performing a joint optimization of all parameters.
Determining the kernel-function form
In this section we derive the functional form of the kernel from conductance relaxation measurements. The decay segments of the conductance evolution are shown in log-log coordinates in Figure˜2, with data that deviate from the linear behavior omitted. It is visible that the initial slopes of the decay are linear up to over two orders of magnitude, which excludes exponential decay. We fit the data using functions of the form $\alpha x+b_{j}$, where $\alpha$ is the common slope and $b_{j}$ is the individual offset.
Figure 2: Log–log representation of post-potentiation conductance relaxation in the PCaPMA memristor. Each curve shows the decay of conductance after a train of trigger pulses with different total stimulation time $T_{j}$. When plotted in double-logarithmic coordinates, the initial segments of all traces are close to straight lines over more than two decades in time, indicating a power-law-like decay and ruling out a simple exponential relaxation.
The function $\alpha x+b$ is transformed into lin-lin coordinates as
$\bar{b}\,\frac{1}{x^{\alpha}}\,$,
for a suitably modified $\bar{b}$. The expression is undefined for $x=0$, so a small positive shift by $\varepsilon$ is introduced. The addition of $\varepsilon$ preserves the short-time characteristic of the relaxation, but avoids the singularity at zero.
Owing to the power-law characteristic, we are considering the kernel function in the form
[…]
for $\alpha>0$. To ensure proper scaling, the kernel must be normalized. To this end set
$I_{\varepsilon}=\int^{\infty}_{\varepsilon}\frac{1}{x^{\alpha+1}}\,dx\,=\frac{1}{\alpha\cdot\varepsilon^{\alpha}},$
and define the resulting normalized kernel as
$\frac{\ker(t)}{I_{\varepsilon}}.$
Finding the kernel-function coefficient $\alpha$
To estimate the value of $\alpha$ that results in the measured conductance $G$, assume the conductance data are the response to a box function of applied voltage $b\,\chi_{(-T,\,0)}$, where $b$ is a scalar and $T$ is the time of applied stimulation.
$G(t)=b\,\chi_{(-T,\,0)}\ast\ker(t)+c\,$,
for a suitable stationary conductance constant $c$. We calculate
$G(t)$ $=b\,\int^{t}_{-\infty}\ker(t-\tau)\,\chi_{(-T,\,0)}(\tau)\,d\tau+c=b\,\int^{0}_{-T}\ker(t-\tau)\,d\tau+c$
$=b\,\int^{T+t}_{t}\ker(u)\,du+c=b\,\big(K(t+T)-K(t)\big)+c$
where
$K(t)=\int^{t}_{-\infty}\ker(\tau)\,d\tau=\begin{dcases}\makebox[44.52858pt][c]{$0\,,$}&\text{if }t<0\\ 1-\dfrac{\varepsilon^{\alpha}}{(t+\varepsilon)^{\alpha}}\,,&\text{if }t\geq 0,\end{dcases}$
is a primitive function to $\ker()$. To derive the decay characteristic after the pulse stimulation from time T-T to 0, fit the decay of the graph to the function for $t\geq 0$
$G(t)=b\,\big(K(t+T)-K(t)\big)+c=b\,\Big(\frac{\varepsilon^{\alpha}}{(t+\varepsilon)^{\alpha}}-\frac{\varepsilon^{\alpha}}{(t+T+\varepsilon)^{\alpha}}\Big)+c$
We repeat the above procedure for every measured trace and find an optimal $\alpha$ over all five datasets, as shown in Figure˜3.
Figure 3: Kernel function coefficient estimation. Each trace corresponds to stimulation by pulses over the duration $T_{j}$. For each trace, a fit $g_{j}(t)=b_{j}\,(\frac{1}{(t+\varepsilon)^{\alpha}}-\frac{1}{(t+\varepsilon+T_{j})^{\alpha}})+c_{j}$ is performed using two trace-specific parameters $b_{j}$ and $c_{j}$ together with the global coefficient $\alpha$ to be estimated.
We have set $\varepsilon$ to the time resolution of the measurements (7070ms) and calculated the value of alpha to be $\alpha=0.029$, which corresponds to the volatility following a power-law with an exponent $1.029$.
Derivation of cumulative conductance
In this section we describe a method for determining the form of the cumulative conductance function $H$. We do so using data from pulse stimulation, again idealized by a single box function similarly as above. The procedure yields only the dependence on the parameter $s$, the dependence on $w$ must be copied, or derived in a different way.
Our goal is to find the cumulative conductance function such that the prediction given by the model results in the traces $g_{j}$ from Figure˜3. Note that because the following utilizes the Laplace transform, we must restrict ourselves to positive times, hence in contrast to the discussion above we work with the single box function supported on the interval (0,T)(0,\,T), i.e. $\chi_{(0,\,T)}$.
  1. 1.
    We use the previously calculated traces dependent on $T_{j}$:
    $g_{j}=b_{j}\,\ker\ast\chi_{(0,\,T)}+c_{j}\,.$
  2. 2.
    Find explicit functions $b(T)$ and $c(T)$ that describe the dependence of $b_{j}$ and $c_{j}$ on $T_{j}$, respectively. In our case, we find by minimization appropriate constants $b^{i}$, $c^{i}$ such that
    $b(T)$ $=b^{1}\,\log(T+b^{2})+b^{3}$
    $c(T)$ $=c^{1}\,\log(T+c^{2})+c^{3}.$
    Then the traces become
    $g=b(T)\,\ker\ast\chi_{(0,\,T)}+c(T)\,.$ (10)
  3. 3.
    Neglecting the saturation term, our model predicts $g=\ker\ast dH(s)+g_{0}$, where $dH(s(t))=\partial_{s}H(s(t))\,\dot{s}(t)$. Recall the voltage-control linear differential equation Eq.˜2, $\dot{s}=\mu\,v$, and consider a normalized voltage input $v=\chi_{(0,T)}$. Then it holds that $s(t)=\mu\,t$ and by setting Eq.˜10 equal to our model we obtain:
    $\ker\ast\big(\partial_{s}H(s)\,\mu\,\chi_{(0,T)}\big)\,+g_{0}=b(T)\,\ker\ast\chi_{(0,\,T)}+c(T)\,.$
  4. 4.
    We apply the Laplace transformation to both sides and perform some rearranging to derive the form of $dH$, valid for $t\in[0,T]$:
    $\mathcal{L}\big(\ker\big)\,\mathcal{L}\big(\partial_{s}H(s)\,\mu\,\chi_{(0,T)}\big)\,+\frac{1}{s}g_{0}=b(T)\,\mathcal{L}\big(\ker\big)\mathcal{L}\big(\chi_{(0,\,T)})+\frac{1}{s}c(T)$
    $\mu\partial_{s}H(\mu\,t)=\mathcal{L}^{-1}\Big\{\,\big(b(T)\,\mathcal{L}(\ker)\,\mathcal{L}(\,\chi_{(0,\,T)})+\frac{1}{s}(c(T)-g_{0})\big)/\mathcal{L}(\ker)\,\Big\}\,.$
  5. 5.
    In our case, the calculation of the Laplace transform of $\ker()$ and $\chi_{(0,T)}$ is performed, up to normalization constant, as
    $\mathcal{L}(\ker)(s)$ $=s^{\alpha}e^{s\varepsilon}\,\Gamma(\alpha,\,s\varepsilon)$
    $\mathcal{L}(\,\chi_{(0,T)})(s)$ $=\frac{1}{s}\,(1-e^{-sT})\,$,
    where $\Gamma(a,x)=\int_{x}^{\infty}t^{a-1}\,e^{-t}\,dt$ is the upper incomplete Gamma function.
    As a result, the above procedure yields the following expression for positive values of $x$:
    $\partial_{s}H(x)=h_{1}\,x^{a}+h_{0}\,.$
    In general there may be some dependence of $H$ on $T$ (refer to Figure˜4), and we choose the parameters $h_{1},\,h_{0}$ of the cumulative function that minimizes such dependence. To obtain $dH$, extend the relationship symmetrically for $x<0$, hence the resulting functional form is
    $dH(s)=(h_{1}\,|s|^{a}+h_{0})\,\dot{s}$ (11)
Figure 4: Cumulative conductance function estimate. Each curve shows the estimated dependence of $\partial_{s}H(s)$ on the state variable $s$, reconstructed from relaxation traces with different stimulation durations $T$ using the discussed procedure. The parameters $h_{0}$, $h_{1}$ are chosen so that the traces for all $T$ collapse onto a power-law-like curve with a common a=0.585a=-0.585, well approximated by the parametric form $dH(s)=(h_{1}\,|s|^{a}+h_{0})\,\dot{s}$. This justifies the simple functional form of dHdH used in the main text.
Finding cumulative conductance function coefficients
Having determined the characteristic of the decay, we now turn to estimating the parameters for the remaining parts of the model. Neglecting the saturation component for now, the resulting conductance $G$ is predicted by the model as:
$G=\ker\ast dH+g_{0}\,.$
We are considering $dH$ in the form derived in Eq.˜11:
$dH(s)=(h_{1}\,|s|^{a}+h_{0})\,\dot{s},$
where we already have estimated the optimal coefficient a=0.585a=-0.585, and are looking to estimate the parameters $h_{0}$ and $h_{1}$. For demonstration, we perform the fit on the stimulation segment of the traces above, this time with the precise input data. We obtain the graphs in Figure˜5.
Figure 5: Comparison between the model and the experimental data for stimulation input of $T=35$ s and $T=3.5$ s. The measured conductance decays after potentiation by a pulse train of total duration $T$ and the solid line shows the corresponding numerical simulation of the model. The overall good agreement across the entire decay window indicates that a single, material-specific kernel captures the volatile dynamics for different stimulation protocols.
Modeling STDP
To estimate the coefficients of the STDP component, we extend the previously computed part of the model with the dependence on the weight parameter $w$. Specifically we are now looking to estimate the coefficients from Eq.˜3: $A_{\mathrm{+}},\,A_{\mathrm{-}},\,\tau_{\mathrm{+}},\,\tau_{\mathrm{-}}\,$, as well as the coefficients from the STDP component of $dH$; $h_{01}$ and $h_{00}$.
The data under evaluation consists of three conductance measurements per run. The first determines the steady-state conductance $g_{0}$, while the other two quantify the conductance change $\Delta G$: one measured 10ms after the end of the second testing pulse (denoted $t_{1}$) and another measured after 100ms (denoted $t_{2}$). We optimize the parameters for the first measurement, since the second one should be a result of the volatility modeled by $\ker()$.
It should be noted, that this estimate is considered to independent of the previous discussion, since any effect arising from the variable $s$ is effectively canceled out by the use of equal but opposite pulses in these measurements and simulations.
We calculate the difference in resting and STDP-induced conductance
$\Delta G=\frac{\ker\ast dH\,(t_{1})-g_{0}}{g_{0}}\,$,
and estimate the parameters of the model to fit the measurements. The results of the STDP-induced conductance at both sample times are shown in fig.˜6. As a consequence, we verified that the decrease in STDP effects over time is a result of the previously modeled volatility mechanism.
Figure 6: Fit and prediction of the STDP-like behavior. The dots show the experimentally measured relative conductance changes $\Delta G/G_{0}$ as a function of spike-timing difference $\Delta t$ for a paired-pulse protocol applied to the PCaPMA memristor. The black lines represent the model output with parameter values fitted to the data at reading delay $t_{1}=10$ ms. The red lines represent the prediction of the model, with the same parameters, when evaluated at reading delay $t_{2}=100$ ms.
Implementing saturation
Up to this point, we have ignored the effects of saturation in the modeling for simplicity of presentation. We thus implicitly assumed the effects modeled so far lie in the linear regime of the saturation curve, so that the subsequent addition of the saturation component will have minimal effect.
The above discussed approach may be repeated with chosen coefficients for the saturation module to obtain optimized model parameters that now include the effects of saturation. Suitable coefficients for $g_{\mathrm{max}},\,g_{\mathrm{min}},\,h_{\mathrm{max}}$ and $h_{\mathrm{min}}$ are presented in the Appendix.
Methods
Experimental data
The proposed model is validated against experimental data obtained from polymeric memristive devices that were synthesised and experimentally characterized by Panthi et al. [16]. In low voltage regime, these devices exhibit a rich set of synaptic-like behaviors, including history-dependent conductance changes and volatile memory, which provide the experimental grounding for our model’s structure.
Polymeric memristor with neuro-synaptic properties
The experimental validation of our model employs memristive devices based on poly[N-(3-(9H-carbazol-9-yl)propyl)methacrylamide] (PCaPMA). In these devices, a thin PCaPMA film is sandwiched between an indium tin oxide (ITO) bottom electrode and a top electrode made of Au or Al. The resulting structure functions as a one-port memristive element whose conductance depends on the history of applied voltage pulses.
Panthi et al. demonstrated that such PCaPMA devices can emulate a broad repertoire of synaptic functions that are central for neuromorphic computing [16]. When excited by trains of low-amplitude voltage pulses, the device conductance can be gradually increased (potentiation) or decreased (depression), depending on the polarity, number, amplitude, and frequency of the pulses. The same platform supports both short- and long-term plasticity, paired-pulse facilitation and depression (PPF/D), spike-timing dependent plasticity (STDP), and simple associative learning protocols, all on millisecond timescales. These characteristics make the PCaPMA memristor a natural candidate for validating compact models that aim to unify memristive switching, volatility, and synaptic plasticity.
The microscopic origin of the resistive switching in PCaPMA has been attributed to a combination of voltage-induced conformational changes of the polymer, trapping and detrapping of charge carriers at localized sites, and redox phenomena within the active layer [17, 16]. A detailed microscopic description of these intertwined processes is beyond the scope of the present work; instead, we treat them as defining an effective internal state whose dynamics is captured phenomenologically by the state variables of our model and by the cumulative conductance function $H(s,w)$.
Beyond the synaptic functions reported in Panthi et al. [16], our collaboration with the device-fabrication team provided access to extensive relaxation and timing datasets obtained on the same class of devices, including unpublished measurements. These data show that, after potentiation by a finite pulse train, the conductance relaxes slowly over several decades in time, with decay curves that are close to linear when plotted in double-logarithmic coordinates (Figure˜3). This long-tailed relaxation is inconsistent with a single exponential and forms the empirical basis for our choice of a power-law-like memory kernel in the volatility module.
For modeling purposes, the key device-level properties that we retain from the PCaPMA system are therefore: (i) analog, gradual conductance modulation by trains of voltage pulses; (ii) coexistence of volatile and non-volatile plasticity on millisecond–second timescales; and (iii) low-voltage operation compatible with mixed-signal neuromorphic circuits. While our framework is formally material-agnostic, PCaPMA provides a concrete polymer platform whose experimentally characterized synaptic-like behavior and volatile relaxation impose quantitative constraints on the kernel $\mathrm{ker}(t)$, the cumulative conductance $H(s,w)$, and the saturation module.
Discussion
We have introduced a modular memristor model that unifies a voltage-controlled memristive core, synaptic-like plasticity and volatile relaxation within a single, computationally efficient framework. By grounding the modules for volatility and plasticity in principles from condensed matter physics and computational neuroscience, respectively, the model moves beyond pure phenomenology to provide a physically and biologically plausible description of complex memristive behaviors.
Extending the state-of-the-art memristor models
The current landscape of compact memristor modeling is dominated by frameworks such as the TEAM and VTEAM models, which provide computationally efficient descriptions of the core resistive switching but are fundamentally phenomenological [13, 14]. Other approaches have examined the dynamical aspects of memristive switching, including shock-wave propagation phenomena that govern commutation speed [24]. Our model extends this state-of-the-art by incorporating two crucial, biologically inspired features—synaptic plasticity and volatile memory—not as ad-hoc additions, but as modules derived from established theoretical principles.
In neuromorphic contexts, the STDP module enables biologically plausible learning rules, facilitating simulations of associative memory and adaptation [10, 29, 20].
To model synaptic plasticity, we implement a spike-timing-dependent plasticity (STDP) rule inspired by the work of Gerstner et al. [6]. This approach is based on the concept of eligibility traces, a cornerstone of modern theories of learning in the brain. In this framework, plasticity is a two-stage process. First, the near-coincidence of pre- and post-synaptic activity creates a short-lived synaptic "tag" or trace that marks the synapse as "eligible" for modification. Second, this temporary trace is only converted into a lasting weight change if a third, modulatory signal—representing global factors like reward, novelty, or surprise—arrives while the trace is still active [7]. By adopting this mechanism, our model moves beyond simple pairwise STDP to be extended toward and flexible three-factor learning rules. This capability aligns the model with cutting-edge research in reinforcement learning and makes it a compelling component for building truly adaptive neuromorphic systems.
To model memory volatility, we took inspiration from the theory of linear viscoelasticity, which describes history-dependent phenomena in materials science [15]. The mathematical structure is that of a hereditary integral, where the current conductance state is a convolution of the entire past history of voltage stimuli with a decaying memory kernel. On top of this linear convolution, we add a final nonlinearity to shape the output conductance. This structure is analogous to the well-known linear-nonlinear models widely used in computational neuroscience to describe the relationship between a stimulus and a neuron’s firing rate [6]. In our model, the linear filter captures the device’s history-dependent memory, while the subsequent nonlinear stage accounts for saturation and other device-specific characteristics.
In contrast to the V-VTEAM model [18], which implements a stretched-exponential volatility model directly in the formulation of the dynamics, our approach to volatility is threshold independent and offers a broad range of possible decay characteristics. For example, a differential equation formulation for a $1/t$ kernel is notoriously computationally unstable, but poses no difficulty in the convolution approach.
Physical origin of the 1/t-like memory kernel
The volatility module of our model is governed by the hereditary kernel in Eq.˜9 with an exponent $\alpha\approx 0.03$ obtained from the log–log relaxation data in Figure˜2. The corresponding cumulative memory $K(t)=\int_{-\infty}^{t}\ker(\tau)\,d\tau$ decays approximately as $K(t)\sim 1/t$ for large $t$. In practice, the device therefore exhibits effectively $1/t$-like “loss of memory” over time.
Power-law relaxations of this type are typical for transport in disordered media. In a broad class of disordered conductors, charge transport proceeds along percolating networks of microscopic pathways, and charge carriers perform a continuous-time random walk (CTRW) with a broad distribution of waiting times between hops [12, 19, 22]. If the waiting-time distribution $\psi(\tau)$ has a heavy tail, e.g. $\psi(\tau)\sim\tau^{-1-\alpha}$ with $0<\alpha<1$, the resulting transport is sub-diffusive and the macroscopic response to a perturbation relaxes as a power law rather than as a sum of exponentials. In our context, the “response” is the conductance of the memristive device following a potentiating pulse train, and the observed long-tailed decay directly motivates the $1/t$-like kernel in Eq.˜9.
The viscoelastic analogy provides a convenient heuristic explanation of this situation. In linear viscoelasticity, a material with a broad spectrum of relaxation times $\rho(\tau)$ can be represented as a superposition of many Maxwell modes, and a spectrum $\rho(\tau)\propto\tau^{-1-\alpha}$ yields an effective relaxation modulus $G(t)\propto t^{-\alpha}$ [15, 21]. Our choice of $\ker(t)\propto(t+\varepsilon)^{-(\alpha+1)}$ plays the same role: it compactly encodes the presence of many internal relaxation processes without having to resolve them individually at the level of the model.
For the PCaPMA devices used here, the microscopic origin of such a broad relaxation spectrum is likely associated with structural and energetic disorder in the polymer, together with charge trapping/detrapping and redox processes in the active layer, as discussed in Panthi et al. [16]. Different local environments and slow morphological rearrangements would naturally generate a hierarchy of timescales. Our model does not attempt to distinguish these mechanisms explicitly. Instead, the 1/t-like kernel should be viewed as a phenomenological summary of their cumulative effect, constrained by the experimentally observed long-tailed relaxation in Figures˜2 and 3.
Importantly, the argument leading to a power-law kernel is not specific to PCaPMA. Percolation-controlled transport and CTRW-style dynamics are common in organic and hybrid memristive systems [25]. We therefore expect that, when adapting the present framework to other polymeric or soft-matter memristors, only the kernel exponent $\alpha$ and an overall time scale need to be adjusted, while the basic structure of the volatility module – a viscoelastic-type kernel acting on the cumulative conductance function $H(s,w)$ – remains unchanged.
Adapting model to polymeric memristor
[…]
Prospective model extensions
The modularity of the proposed model is one of its key strengths and opens several avenues for future work. Each component—the non-volatile core, the volatility kernel, and the plasticity rule—can be modified or replaced independently to adapt the model to different physical devices or to explore alternative theoretical assumptions. This allows for the separate experimental characterization and fitting of each module, simplifying the process of model validation and refinement for new memristive technologies.
Furthermore, while the current model is deterministic, the experimental data clearly exhibit stochasticity. This inherent randomness in polymeric devices is not necessarily a flaw; it can be a functional feature. For example, recent work has demonstrated that the stochastic switching of polymeric memristors can be used as a physical entropy source for applications in probabilistic computing [4]. A natural extension of our work would be to incorporate a stochastic component into the model, which would not only improve its descriptive accuracy but also allow it to be used for designing and simulating such novel computing paradigms.
The $1/t$ form is phenomenological; multiple microscopic routes (trapping/detrapping spectra, filament reconfiguration, ionic drift) can lead to similar envelopes. Future work should correlate $\alpha$ with material properties (electrode, thickness, processing) and examine frequency dependence.
Conclusions
We have presented a modular and computationally efficient compact model for memristors that exhibits both volatile memory and synaptic-like plasticity. The model’s key innovation lies in the synthesis of several modeling techniques that allow for memristor current-voltage characteristics, synaptic-like plasticity effects and a general description of relaxation dynamics with saturation. We have discussed how the volatile relaxation dynamics, experimentally observed as a power-law decay, can be described by a viscoelasticity-inspired hereditary integral whose form is physically motivated by percolation theory. Furthermore, we have demonstrated that a biologically plausible model of STDP based on eligibility traces can be implemented, endowing the future devices with a potential capacity for advanced learning.
The complete model was shown to be in excellent quantitative agreement with experimental data from polymeric memristors. Furthermore, we believe that the modularity and relative simplicity of the proposed model allow its implementation (including validation) in a wide range of memristive-like devices. Therefore, this work provides a unified, mathematically grounded framework for understanding and simulating complex memristive devices, offering a valuable tool for advancing the design of neuromorphic computing systems.
Funding
This work was supported by The Czech Science Foundation project No. 24-10384S, MEYS project Inter-Excellence No. LUAUS24032 and the Strategy AV 21 Programme “Breakthrough Technologies for the Future – Sensing, Digitisation, Artificial Intelligence, and Quantum Technologies”.
Author contributions
D.H. and L.K. conceived and led the project, carried out the theoretical development, implemented the model and simulations, analyzed the data, and wrote the manuscript. S.H.F. and K.K. contributed to the theoretical development and provided critical feedback and revisions to specific sections of the manuscript. A.P., Y.R.P., J.P., and J.V. fabricated the polymeric memristors and performed the experimental measurements.
Competing interests
The authors declare no competing interests. All authors reviewed the manuscript and approved the final version.
Data availability
The source data of PPF/PPD and STDP can be accessed from Zenodo at https://zenodo.org/records/12685756.
The source code for the model can be accessed at https://github.com/danielhabart/modular_memristor
Appendix: Numerical values
Here we collect numerical values of the parameters fitted in the preceding sections. They are not required for understanding the structure of the model, but they provide a complete description of the specific PCaPMA device and protocols used in our study.
The first Table˜1 lists the kernel exponent $\alpha$. The second table˜2 summarizes the amplitudes $b$ and offsets $c$ of the relaxation traces for each stimulation duration $T$, which determine the functions $b(T)$ and $c(T)$ appearing in the cumulative conductance function-reconstruction procedure Eq.˜10. The cumulative conductance function power law coefficient is shown in Table˜3. The Table˜4 lists cumulative conductance coefficients $h_{11}$ and $h_{10}$, as well as model coefficients $\mu$ and $g_{0}$, estimated to fit the potentiation segment of Figure˜5.
Table˜5 reports the parameters of the STDP module (Eq.˜3), which are used to generate the model STDP curve in Figure˜6. Our choice for the parameters of the saturation module are presented in Table˜6.
Together, these values define a reproducible parameter set for the particular device studied, while keeping the main text focused on the general modeling framework.
$\alpha$ 0.029
Table 1: Volatility power law coefficient estimate.
$T$ [s] 70 35 7 3.5 0.7
$b$ 54.65 41.44 40.23 33.07 29.78
$c$ 30.31 30.43 30.63 30.98 31.13
Table 2: Numerical values of the trace-specific amplitudes $b$ and offsets $c$ obtained from fitting the relaxation traces for different stimulation durations $T$. These values determine the functions $b(T)$ and $c(T)$ in Eq.˜10, which are used to derive the kernel function $\ker(t)$ and, in turn, the form of the cumulative conductance function $H(s)$ from the experimental data.
$a$ -0.585
Table 3: Cumulative conductance function power law coefficient estimate.
$T$ $h_{11}$ $h_{10}$ $\mu$ $g_{0}$
[…]
Table 4: Estimate of model parameters. Owing to the stochastic nature of the polymer-based memristor under study, we obtain a good fit per measured trace of potentiation stimulation.
$A_{\mathrm{+}}$ 10.9
$A_{\mathrm{-}}$ 0.90
$\tau_{\mathrm{+}}$ 0.176 s
$\tau_{\mathrm{-}}$ 0.619 s
$h_{\mathrm{01}}$ $1.01\times 10^{-4}$
$h_{\mathrm{00}}$ $2.52\times 10^{-3}$
Table 5: Numerical values of the parameters of the STDP module. The parameters $_{\mathrm{+}}, AA_{\mathrm{-}}$ and $\tau_{\mathrm{+}}$, $\tau_{\mathrm{-}}$ define the shapes and time constants of the potentiating and depressing lobes in the local-variable STDP dynamics Eq.˜3, the coefficients $h_{\mathrm{01}}$ and $h_{\mathrm{00}}$ determine the dependence of the cumulative conductance function $H$ on the variable $w$ as in Eq.˜6. These values are used to generate the model STDP curve shown in Figure˜6 and to ensure consistency between the dynamical STDP formulation and the measured memristor response.
[…]
Table 6: The saturation module coefficients were chosen to minimize the effect of added saturation to the model. The parameters $h_{\mathrm{max}},\,h_{\mathrm{min}}$ were set equal to $g_{\mathrm{max}}$, $g_{\mathrm{min}}$, respectively.